OpenStudy (anonymous):

so im in math and i got this question and I have no idea how to Q: Marvel Man movie is being released to a limited number of movie theaters. The number of showtimes for the movie is twice the number of movie theaters. The number of seats available for each showtime is three times the number of showtimes. Let x represent the number of theatres showing the movie. Determine the three factors that are to be multiplied to determine the total number of seats available for the movie. When simplified what is the degree and classification of this expression? Inte

OpenStudy (jordan123321):

what do you think

OpenStudy (anonymous):

Honestly have no idea xD

OpenStudy (anonymous):

Number of theaters showing the movie = x "number of showtimes for the movie is twice the number of movie theaters" Number of showtimes for the movie = 2 * x "number of seats available for each showtime is three times the number of showtimes" 1) Number of seats available for each showtime = 3 * 2 * x = 6x The three factors that need to be multiplied to determine the TOTAL number of seats are: 6x, 2x and x. :D

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

yep :D

OpenStudy (jordan123321):

no direct answers

OpenStudy (anonymous):

well thanks guys xD i feel totally stupid not knowing how to do it, but I guess i should stop napping during math xDDD

OpenStudy (anonymous):

>-> you got it @haleyritchie1223 ?

OpenStudy (anonymous):

ha nice naps are good! XD

OpenStudy (anonymous):

xD i need my sleep sometimes xDDD

OpenStudy (anonymous):

and all they do is talk so xDD

OpenStudy (anonymous):

XD